In a triangle ABC, AB=AC and angle A = 50, find angle B and angle C

Grade:9

3 Answers

Phanindra
6 years ago
If two sides are equal then their opposite angles are also equal...this is isoceles triangle property...
So angle B and C are equal....
So...by angle sum property...
A + B + C = 180...
B + B = 180 - 50... 
             = 130 
Sooo.....B = C = 65
Sugnyan S Abblgeri
6 years ago
Triangle ABC is an isosceles triangle because the two sides are equal that is AB=AC
Angle A =50
So in isosceles triangle the angles opposite to equal sides are equal
So Angle B=Angle C
Then Angle C can be written as Angle B
now by angle sum property
A+B+C=180
A+B+B=180
A+2B=180
50+2B=180
2B=180-50
2B=130
B=130/2
B=65
So Angle C =65
Therefore Angle B=65 and Angle C=65
 
 
Bharat Luthra
5 years ago
In an isosceles triangle, the angles opposite to equal sides are equal.
so angle B=angle C
By angle sum property of triangle,
angle A+angle B+angle C=180 degree
or 50 degree+angle B+ angle B(an angle B=angle C)=180 degree
or 2*angle B=180 – 50 degree
or 2*angle B=130 degree
or angle B=130/2 degree
or angle B=65 degree
thus angle B=65 degree and angle C= 65 degree